(25%) Problem 2: Consider a hydrogen atom with a proton at the center and an electron in the orbit. The proton and the electron are separated by 5.3 x 10-10 meters. The Proton has a charge of 1.6 x 10-19 C, and the electron has a charge of -1.6 x 10-19 C. Take Coulombs constant k to be 9 x 109 N-m?/c2 tud Grad Dedu Potent sin cotan) asin) tan() | π acosO Subm Attem cosO 78 9 detail atan) acotan) sinh(0 coshOtanh0 cotanh0 Degrees O Radians 0 END BACKSPACE CLEAR Submit I give up! Hits: 0 for a 0% deduction . Hints remaining Feedback: 0% deduction per feedback 50% Part (b) What is the magnitude of the electrical force, in newtons, the electron exerts on the proton?
Answer #1

Solution)

Part a) We know , Force F= k*q1*q2/r^2

So, F= (9*10^9)*(1.6*10^-19)/(5.3*10^-10)^2

F= 5.12*10^9 N (Ans)

====

Part b) Here also it will be same.

======

Good luck!:)